# Rampa: Robotic Augmented Reality for Machine Programming and Automation

### v1

@@ Abstract
As robotics continue to enter various sectors beyond traditional industrial applications, the need for intuitive robot training and interaction systems becomes increasingly more important. This paper introduces Robotic Augmented Reality for Machine Programming (Rampa), a system that utilizes the capabilities of state-of-the-art and commercially available AR headsets, e.g., Meta Quest 3, to facilitate the application of Programming from Demonstration (PfD) approaches on industrial robotic arms, such as Universal Robots UR10. Our approach enables in-situ data recording, visualization, and fine-tuning of skill demonstrations directly within the user’s physical environment. Rampa addresses critical challenges of PfD, such as safety concerns, programming barriers, and the inefficiency of collecting demonstrations on the actual hardware. The performance of our system is evaluated against the traditional method of kinesthetic control in teaching three different robotic manipulation tasks and analyzed with quantitative metrics, measuring task performance and completion time, trajectory smoothness, system usability, user experience, and task load using standardized surveys. Our findings indicate a substantial advancement in how robotic tasks are taught and refined, promising improvements in operational safety, efficiency, and user engagement in robotic programming.
@@
augmented reality, extended reality, mixed reality, programming from demonstration
@@ I Introduction
The advent of sophisticated robotic applications across various industries necessitates advanced programming techniques that can accommodate complex, dynamic environments. The conventional approach to robot programming typically involves segmented processes where users must alternate between physical and digital realms. Traditional approaches rely heavily on offline simulation and manual coding, which can be time-consuming and often impractical for real-time applications [1]. As these methods lack the capability to adapt quickly to new environments or to intuitively incorporate human expertise, the complexity and time involved in programming increases, which also heightens the risk of errors and accidents.
The procedure known as Programming from Demonstration (PfD), a.k.a. Learning from Demonstration (LfD), is a widely adopted procedure that enables robots to learn skill policies by watching an expert [2]. Recently, Machine Learning (ML)-based PfD frameworks have become instrumental in enabling robots to model and execute intricate skills autonomously. Yet, integrating such approaches into robotic programming often encounters significant challenges, including the labor-intensive and potentially hazardous processes involved in data gathering and model testing in the real world  [3].
In response to these challenges, we introduce a sophisticated AR-based framework that transforms robotic programming through direct, interactive modeling and visualization. Our system, Robotic Augmented Reality for Machine Programming and Automation (Rampa), leverages the immersive and interactive capabilities of state-of-the-art XR headsets (e.g., Meta Quest 3) to create a cohesive and safe user experience for robot programming and automation of tasks in situ.
Figure 1: With Rampa, users can demonstrate trajectories and simulate robot movements in situ.
Figure 2: Rampa allows users to (a) collect trajectory data by drawing the desired path with a pinch. (b) The virtual robot then simulates the trajectory, allowing users to observe how the actual robot would interact with its environment, given the trajectory. (c) After the simulation, the user may demonstrate the trajectory on the actual robot. (d) Rampa also facilitates the collection of trajectory data to train ML models. After training, users can condition the model interactively and generate a sample trajectory from it, as shown in Fig. 5.
By integrating in-situ simulation, real-time spatial understanding, and direct manipulation capabilities, Rampa offers a dual advantage: it streamlines the programming process while enhancing the physical safety, flexibility of robotic operations, and efficiency of PfD procedures and, therefore, reduces the barriers to effective human-robot collaboration. The ability to produce and adjust demonstration trajectories in real-time and in situ facilitates the task of the operator while interacting directly with the learned models establishes a deeper understanding of these models.
@@
Operators use XR to register real and/or virtual objects and demonstrate skills, capturing the data directly in the user’s environment. Compared to traditional methods, such as the kinesthetic control, Rampa simplifies the collection of demonstration trajectories, an integral component of PfD procedures, offering a more efficient experience for robotic programmers.
@@
Utilizing the collected data, popular PfD models, such as Probabilistic Movement Primitives (ProMPs) [4], can be trained and validated within the AR environment. Operators can observe the generated trajectories in real-time, ensuring the robot’s behavior aligns with desired outcomes.
To the best of our knowledge, we propose the first end-to-end XR-driven PfD framework to enable the demonstration of robotic trajectories, training, and testing ML model augmented by visualization of the robot motion and trajectory adjustment in situ. Importantly, our comprehensive evaluation comprises the experiment consisting of three distinct robotic tasks, 20 participants including both experienced and novice users, and an extensive analysis using quantitative metrics, incorporating standard surveys, such as System Usability Scale (SUS) [5], User Experience Questionnaire (UEQ) [6], and NASA Task Load Index (NASA-TLX) [7].
@@
  • The release of the system as open-source for the robotics community, compatible with cutting-edge mixed reality headsets (e.g., Quest 3) and robotic arms.
  • Comprehensive evaluation and analysis of the system in teaching different robot manipulation tasks using quantitative metrics measuring task performance and completion time, trajectory smoothness, system usability, user experience, and task load.
  • Facilitation of interactive and situated programming of PfD models suitable for commonly used movement primitives in robotics.
@@ II-A Movement Primitives in Programming from Demonstration
Enabling robots to perform intricate real-world tasks has been the driving force of robotics research. Functioning in the real world, outside the controlled laboratory conditions, requires sophisticated policies. To overcome the difficulty of developing such policies, movement primitives (MP), which provide effective abstractions that facilitate motor control [8], is a modular formalism widely employed to encode and reproduce complex movements. PfD enables learning MP by observing an expert [2]. Following the PfD paradigm, [8] proposed the Dynamical Movement Primitives (DMPs) model to learn MPs by leveraging human expertise. The original DMP model could encode only a single trajectory, while different demonstrations may contain important variabilities that the modeling framework must capture [9, 10]. In order to capture the variety in demonstrations, [11] proposed to model movement trajectories using probabilistic methods such as GMM and GMR, and [12] combined the DMP and GMM based approaches. These methods allowed the use of multiple demonstrations to model the skill trajectories, enhancing the representational power and the generalization capacity of MPs [13].
Latest advances in ML have significantly influenced PfD approaches, enabling the synthesis of more flexible and generalizable MP-based controllers to perform tasks with higher efficiency and adaptability. Recent studies, such as [14, 15, 16], use deep neural networks to model movement trajectories in the form of MPs. Although they are successful in encoding and reproducing more complex motor trajectories, they are not practical for real-time applications, specifically for human-in-the-loop interactive systems, as they need a training phase to model demonstration trajectories. Due to the real-time requirement, a widely used approach, ProMPs, which models a distribution of trajectories as a probability distribution over a set of basis functions, improving the generalization capability to novel via points [17], is used as the MP modeling framework in this study.
Despite these techniques allowing for efficient learning and generalization of movements, the application of these methods generally requires time-consuming procedures in data collection, model training, and applying the trained model back to real-world settings [3]. Furthermore, operators could find it challenging to grasp and interpret the behavior of ML models, posing challenges to the explainability of AI in robotics. In the Rampa system, integrating AR into the ML training and visualization process allows users to define both real and virtual objects and demonstrate tasks within an AR environment, eliminating manual spatial data recording. Moreover, the Rampa system enhances the comprehensibility of ML models by offering a user-friendly and interactive platform for ML training and fine-tuning.
@@ II-B AR-Oriented Interfaces
As robotic systems have increasingly been employed in various domains, the necessity for efficient methods and intuitive interfaces for robot programming has grown to assist users ranging from novice to technical expert. Different types of robot programming frameworks, such as visual programming methods  [18] combined with speech-based programming  [19] have been proposed for users to control and instruct robotic systems and aim to allow users to communicate instructions and feedback to robots with ease and high precision, which improves the functionality and usability of robotic systems and enables efficient human-robot interaction. Recently, foundation models have been deployed in real-world robotic applications; nevertheless, they are still inadequate for tasks requiring human involvement [20].
@@
Unlike the traditional methods, XR/AR technologies, with their real-time and contextual feedback directly within the user’s field of view, make it possible to have more natural and efficient interactions between human users and robots. As a subset of XR technologies, they enable the superimposition of computer graphics onto the physical world [21] and have been increasingly utilized to enhance human-robot interaction (HRI), bridging the gap between virtual and real environments. Recent surveys  [22, 3] demonstrate the potential of AR to transform robot programming, providing a detailed taxonomy of AR applications in robotics, including beneficial aspects of AR for robot control, programming, and visualization.
@@
AR visualization of the robot’s motion has been used to simplify the process of planning, observing, and modifying the robot’s trajectory for users. For instance, GhostAR [23] enables users to visualize and adjust robot trajectories in real-time, which improves the accuracy of robot actions in human-robot collaborative operations. Similarly, Quintero et al.  [24] introduces a method that allows users to interact with robot trajectories with ease by using an AR interface to visualize the motion of the robot. Other studies [25, 26] show that visualizing robot movement improves task efficiency and enhances user understanding of robot intentions, hence facilitating human-robot interactions.
@@
AR interfaces have also been used for robot training methods such as PfD and teleoperation, as they offer interactive and intuitive ways to teach robots new skills. Holo-Dex [27] provides users with a broad AR-based framework for robot learning with teleoperation. Alongside, Luebbers et al.  [28] offers in-situ visualizations of robot trajectories with their constraints, and Diehl et al. [29] proposes a system for visualization of learned skills for verification and error detection, therefore providing an intuitive platform for users to convey proper demonstrations to the robot. However, these studies require kinesthetically demonstrated trajectories. Lotsaris et al. [30] introduces an AR system to demonstrate robot trajectories by placing the virtual gripper in the target position. This work, however, does not allow the user to demonstrate the precise path. By bridging the gap left by AR-based interfaces for PfD, we propose a comprehensive, end-to-end system that allows users to demonstrate robotic trajectory, visualize the robot motion, modify the trajectory, train, and test the ML model in situ, hence making robotic programming more accessible and practical.
@@ III Rampa: An End-to-End System for Robotic AR
The Rampa system incorporates ROS server, Unity game engine, and Quest 3 XR headset to provide an intuitive platform for interactive demonstration collection, ML training, and visualization. In this section, features and the implementation of the proposed system are discussed.
@@ III-A System Overview
Figure 3: When the auto-transform option is selected, the virtual robot teleports based on an offset relative to the left controller. When the controller is attached to the robot’s base, the virtual robot aligns with the actual robot.
@@ III-A1 Simulated Robot and Auto-calibration
A simulated robot is present to model the actual robot’s behavior, allowing for accurate testing and validation of trajectory plans. An auto-calibration feature is included, which uses the controller from the headset and a tool that can hold the controller for the virtual robot to be aligned with the actual robot, as shown in Fig. 3. The simulation environment ensures that any trajectory can be tested thoroughly for safety before being deployed on the real robot, also making sure the actual robot will replicate both the location and the motion of the simulated one.
@@ III-A2 Trajectory Playback and Adjustment
Trajectory playback and adjustment are implemented to allow the user to fine-tune the recorded trajectory. The user can go backward or forward, play the rest of the re-winded trajectory, and pause when desired (Fig. 4). If the user decides to discard a portion of the trajectory, they can redraw the trajectory from the current way-point. This interactive approach enables precise control of the robot’s movements and ensures that the final trajectory is both accurate and safe for the desired task.
Figure 4: Users can go back and forth in the trajectory and then can redraw it from the current way-point using the UI menu provided.
@@ III-A4 ML Training and Preview
The recorded trajectories can be either discarded or added to a training set, which can then be viewed or deleted, as shown in Fig. 5. The Rampa allows the user to add demonstrated trajectory to the training set after the simulation of the robot motion and/or physical robot execution. Moreover, the user can modify the recorded trajectory post-simulation and/or post-execution, which enhances the adaptability of the system and flexibility of the user during the demonstration collection process. The training set can then be used to train a model. To test the model, the user can add way-points to condition the trained model using virtual objects, as demonstrated in the right-hand side of Fig. 5. This process facilitates the tasks of training, conditioning, and sampling trajectories efficiently in an AR environment.
Figure 5: Users can save and view the trajectories they have recorded. Users can then train an ML model, condition it using virtual cubes, and generate a new trajectory.

### v2

# Rampa: Robotic Augmented Reality for Machine Programming by DemonstrAtion

@@ Abstract
This paper introduces Robotic Augmented Reality for Machine Programming by Demonstration (Rampa), the first ML-integrated, XR-driven end-to-end robotic system, allowing training and deployment of ML models such as ProMPs on the fly, and utilizing the capabilities of state-of-the-art and commercially available AR headsets, e.g., Meta Quest 3, to facilitate the application of Programming by Demonstration (PbD) approaches on industrial robotic arms, e.g., Universal Robots UR10. Our approach enables in-situ data recording, visualization, and fine-tuning of skill demonstrations directly within the user’s physical environment. Rampa addresses critical challenges of PbD, such as safety concerns, programming barriers, and the inefficiency of collecting demonstrations on the actual hardware. The performance of our system is evaluated against the traditional method of kinesthetic control in teaching three different robotic manipulation tasks and analyzed with quantitative metrics, measuring task performance and completion time, trajectory smoothness, system usability, user experience, and task load using standardized surveys. Our findings indicate a substantial advancement in how robotic tasks are taught and refined, promising improvements in operational safety, efficiency, and user engagement in robotic programming.
@@
Virtual Reality and Interfaces, Learning from Demonstration, Augmented Reality (AR), Extended Reality (XR)
@@ I Introduction
The advent of sophisticated robotic applications necessitates advanced programming techniques that can accommodate complex, dynamic environments. The conventional approach to robot programming typically involves segmented processes where users must alternate between physical and digital realms. Traditional approaches rely heavily on offline simulation and manual coding, which can be time-consuming and often impractical for real-time applications [1]. As these methods lack the capability to adapt quickly to new environments or to intuitively incorporate human expertise, the complexity and time involved in programming increases, which also increases the risk of errors and accidents.
The procedure known as Programming by Demonstration (PbD), a.k.a. Learning from Demonstration (LfD), is a widely adopted procedure that enables robots to learn skill policies by watching an expert [2]. Recently, Machine Learning (ML)-based PbD frameworks have become instrumental in enabling robots to model and execute intricate skills autonomously. Yet, integrating such approaches into robotic programming often encounters significant challenges, including the labor-intensive and potentially hazardous processes involved in data gathering and model testing in the real world  [3].
In response to these challenges, we introduce a sophisticated AR-based framework that transforms robotic programming through direct, interactive modeling and visualization. Our system, Robotic Augmented Reality for Machine Programming by Demonstration (Rampa), leverages the immersive and interactive capabilities of state-of-the-art XR headsets (e.g., Meta Quest 3) to create a cohesive and safe user experience for robot programming and automation of tasks in situ. As shown in Fig. 1, users can directly demonstrate robot trajectories and visualize robot movements while maintaining the awareness of the actual workspace. We also release Rampa as open source for the robotics community, offering compatibility with cutting-edge mixed reality headsets and robotic arms.
Figure 1: With Rampa, novice and expert users can demonstrate trajectories and simulate robot movements in situ using augmented reality (AR).
By integrating in-situ simulation and direct manipulation capabilities, Rampa offers a dual advantage: it streamlines the programming process while enhancing the physical safety, flexibility of robotic operations, and efficiency of PbD procedures and, therefore, reduces the barriers to effective human-robot collaboration. The ability to produce and adjust demonstration trajectories in real-time and in situ facilitates the task of the operator while interacting directly with the learned models establishes a deeper understanding of these models.
@@
Operators use XR to demonstrate skills, capturing the data directly in the user’s environment. Compared to traditional methods, such as kinesthetic control, Rampa simplifies the collection of demonstration trajectories, an integral component of PbD procedures, offering a more efficient experience for robotic programmers.
Figure 2: Rampa allows users to (a) draw trajectories with their bare hand, (b) simulate and observe their effect in situ, and (c) execute them on the real robot. (d) Rampa facilitates the direct and streamlined collection of trajectory data to train ML models. Please refer to text for use cases of the system.
@@
Utilizing the collected data, popular PbD models, such as Probabilistic Movement Primitives (ProMPs) [4], can be trained and validated within the AR environment. Operators can observe the generated trajectories in real-time, ensuring the robot’s behavior aligns with desired outcomes.
To the best of our knowledge, we propose the first end-to-end XR-driven PbD framework to enable the demonstration of robotic trajectories, training, and testing ML model augmented by visualization of the robot motion and trajectory adjustment in situ. An overview of the proposed system can be seen in Fig. 2. Importantly, our comprehensive evaluation comprises the experiment consisting of three distinct robotic tasks, 20 participants including both experienced and novice users, and an extensive analysis using quantitative metrics, incorporating standard surveys, such as System Usability Scale (SUS) [5], User Experience Questionnaire (UEQ) [6], and NASA Task Load Index (NASA-TLX) [7], and qualitative user feedback.
@@
  • Comprehensive evaluation and analysis of the system in teaching different robot manipulation tasks using both qualitative feedback and quantitative metrics measuring task performance and completion time, trajectory smoothness, system usability, user experience, and task load.
  • Facilitation of interactive and situated programming of PbD models suitable for commonly used movement primitives in robotics.
@@ II-A Movement Primitives in Programming from Demonstration
Movement primitives (MP), which provide effective abstractions that facilitate motor control [8], is a modular formalism widely employed to encode and reproduce complex movements. Schaal et al. [8] proposed the Dynamical Movement Primitives (DMPs) model to learn MPs by leveraging human expertise. The original DMP model could encode only a single trajectory, while different demonstrations may contain important variabilities that the modeling framework must capture [9, 10]. In order to capture the variety in demonstrations, Gribovskaya et al. [11] proposed to model movement trajectories using probabilistic methods. These methods allowed the use of multiple demonstrations to model the skill trajectories, enhancing the representational power and the generalization capacity of MPs [12].
Latest advances in ML have significantly influenced PbD approaches, enabling the synthesis of more flexible and generalizable MP-based controllers to perform tasks with higher efficiency and adaptability. Recent studies, such as Pérez-Dattari et al. [13], Blessing et al. [14], Yildirim et al. [15], use deep neural networks to model movement trajectories in the form of MPs. However, they are not practical for real-time applications, specifically for human-in-the-loop interactive systems, as they need a training phase to model demonstration trajectories. Due to the real-time requirement, ProMPs, which models a distribution of trajectories as a probability distribution over a set of basis functions, improving the generalization capability to novel via points [16], is used as the MP modeling framework in this study.
In Rampa, integrating AR into the ML training and visualization process allows users to demonstrate tasks within an AR environment, eliminating manual spatial data recording. Moreover, Rampa enhances the comprehensibility of ML models by offering a interactive platform for ML training and fine-tuning.
@@ II-B AR-Oriented Interfaces
Different types of robot programming frameworks, such as visual programming methods Paxton et al.  [17] combined with speech-based programming Alexandrova et al. [18] have been proposed to allow users to communicate instructions and feedback to robots with ease and high precision, which improves the functionality and usability of robotic systems and enables efficient human-robot interaction. Recently, foundation models have been deployed in real-world robotic applications; nevertheless, they are still inadequate for tasks requiring human involvement [19].
@@
Unlike the traditional methods, AR technologies, with their real-time and contextual feedback directly within the user’s field of view, make it possible to have more natural and efficient human-robot interactions. As a subset of XR technologies, they enable the superimposition of computer graphics onto the physical world [20] and have been increasingly utilized to enhance human-robot interaction (HRI), bridging the gap between virtual and real environments. Recent surveys Walker et al. [21] and Suzuki et al. [3] demonstrate the potential of AR to transform robot programming, providing a detailed taxonomy of AR applications in robotics, including beneficial aspects of AR for robot control, programming, and visualization. Specifically, Jiang et al. [22] compare and contrast multiple interaction interfaces for AR-based robot programming, highlighting how different interface designs impact data collection efficiency and overall user experience.
@@
AR visualization of the robot’s motion has been used to simplify the process of planning, observing, and modifying the robot’s trajectory for users. For instance, Cao et al. [23] enables users to visualize and adjust robot trajectories in real-time, which improves the accuracy of robot actions in human-robot collaborative operations. Similarly, Quintero et al.  [24] introduces a method that allows users to interact with robot trajectories with ease by using an AR interface to visualize the motion of the robot. Walker et al. [25], Rosen et al. [26] show that visualizing robot movement improves task efficiency and enhances user understanding of robot intentions, hence facilitating human-robot interactions.
@@
AR interfaces have also been used for robot training methods such as PbD and teleoperation, as they offer interactive and intuitive ways to teach robots new skills. Arunachalam et al. [27] provides users with a broad AR-based framework for robot learning with teleoperation. Alongside, Luebbers et al.  [28] offers in-situ visualizations of robot trajectories with their constraints, and Diehl et al. [29] proposes a system for visualization of learned skills for verification and error detection, therefore providing an intuitive platform for users to convey proper demonstrations to the robot. However, these studies require kinesthetically demonstrated trajectories. Lotsaris et al. [30] introduces an AR system to demonstrate robot trajectories by placing the virtual gripper in the target position. This work, however, does not allow the user to demonstrate the precise path. Recent works Chen et al. [31] and Chen et al. [32] show the potential of AR interfaces, focusing on the demonstration collection phase. However, they do not offer an end-to-end solution. By bridging the gap left by AR-based interfaces for PbD, we propose a comprehensive, end-to-end system that allows users to demonstrate robotic trajectory, visualize the robot motion, modify the trajectory, train, and test the ML model in situ, hence making robotic programming more accessible and practical.
A comprehensive comparison with existing AR-based systems is provided in Table I, evaluating several key features. The End-to-End feature corresponds to whether the proposed system supports demonstration collection, the training and deployment of the training result on the fly, within a single system. ML in AR indicates whether any ML component is integrated into the AR system. Full Trajectory Modification refers to whether the system has rewinding, redrawing capabilities of the collected trajectory pre- or post-demonstration by the actual robot. Online Hand Mimicking by Robot corresponds to whether the system enables online hand mimicking during demonstration collection by either the simulated or the actual robot. Trajectory Visualization represents the capability to visualize the collected trajectory within AR, where Partial corresponds to constraint visualization of the learned skill in Luebbers et al.  [28], and end-effector visualization in Lotsaris et al.  [30]. Lastly, Trajectory Drawing describes the method of trajectory collection within AR.
TABLE I: Comparison of AR-Based Systems
Feature / Method [24] [28] [30] [31] [32] Rampa
End-to-End No No No No No Yes
ML (Training + Deployment) in AR No Yes No No No Yes
Full Trajectory Modification No No No No No Yes
Online Hand Mimicking by Robot No No No Yes Yes Yes
Trajectory Visualization Yes Partial Partial No Yes Yes
Trajectory Drawing Method Point-based None Point-based Hand gestures Hand gestures Hand gestures
@@ III Rampa: An End-to-End System for Robotic AR
The Rampa system integrates a ROS server, Unity game engine, and Quest 3 XR headset to provide an intuitive platform for interactive demonstration collection, ML training, and visualization. Its interface, implemented in Unity, include a simulated robot, a UI menu, and interactive tools such as virtual markers for model conditioning. Users control the system via hand gestures and controller inputs, leveraging features like trajectory adjustment, auto-calibration, online hand following, and ML integration.
@@ III-A1 Simulated Robot and Auto-Calibration
A simulated robot is present to model the actual robot’s behavior, allowing for accurate testing and validation of trajectory plans. For deployment of demonstrations to actual robot, the simulated and the actual robot should be precisely aligned. This problem is addressed by the auto-calibration feature. The feature uses the controller from the headset and a 3D-printed tool that can hold the controller for the virtual robot to be aligned with the actual robot, as shown in Fig. 2-A. The simulation environment ensures that any trajectory can be tested thoroughly for safety before being deployed on the real robot, also making sure the actual robot will replicate both the location and the motion of the simulated one.
@@ III-A2 Trajectory Playback and Adjustment
Trajectory playback and adjustment are implemented to allow the user to fine-tune the recorded trajectory. When a user decides to inspect the current trajectory with the simulated robot, before or after real robot demonstration, they can go backward or forward in the trajectory, play the rest of the re-winded trajectory, and pause when desired (Fig. 3), using the UI provided in the system. If the user decides to discard a portion of the trajectory, they can redraw the trajectory from the current way-point. This interactive approach enables precise control of the robot’s movements and ensures that the final trajectory is both accurate and safe for the desired task.
Figure 3: The user (a) draws a trajectory to the hoop, (b) rewinds the trajectory and chooses to redraw it from the current way-point using the UI menu, (c) redraws the trajectory, and (d) finalizes the trajectory.
@@ III-A3 Real-Time Hand Following
Figure 4: Users can save and view trajectories they have recorded, then train an ML model, condition it using virtual cubes, and generate a new trajectory.
@@ III-A4 ML Training and Preview
The recorded trajectories can be either discarded or added to a training set, which can then be viewed or deleted, as shown in Fig. 4. Rampa allows the user to add demonstrated trajectory to the training set after the simulation of the robot motion and/or physical robot execution. Moreover, the user can modify the recorded trajectory post-simulation and/or post-execution, which enhances the adaptability of the system and flexibility of the user during the demonstration collection process. The training set can then be used to train a model. To test the model, the user can add way-points to condition the trained model using virtual cubes, as demonstrated in Fig. 4. This process facilitates the tasks of training, conditioning, and sampling trajectories efficiently.